OpenStudy (anonymous):

select the property that allows the statement 3=x to be written x=3

OpenStudy (mathstudent55):

Identity for addition: a + 0 = a Commutative: a + b = b + a Associative: (a + b) + c = a + (b + c) Symmetric: a = b means b = a Reflexive: a = a